# Explainable and externally validated machine learning for neuropsychiatric diagnosis via electrocardiograms

### v1

@@ Abstract
Background: Electrocardiogram (ECG) analysis has emerged as a promising tool for identifying physiological changes associated with neuropsychiatric conditions. The relationship between cardiovascular health and neuropsychiatric disorders suggests that ECG abnormalities could serve as valuable biomarkers for more efficient detection, therapy monitoring, and risk stratification. However, the potential of the ECG to accurately distinguish neuropsychiatric conditions, particularly among diverse patient populations, remains underexplored.
Methods: This study utilized ECG markers and basic demographic data to predict neuropsychiatric conditions using machine learning models, with targets defined through ICD-10 codes. Both internal and external validation were performed using the MIMIC-IV and ECG-View datasets respectively. Performance was assessed using AUROC scores. To enhance model interpretability, Shapley values were applied to provide insights into the contributions of individual ECG features to the predictions.
Results: Significant predictive performance was observed for conditions within the neurological and psychiatric groups. For the neurological group, Alzheimer’s disease (G30) achieved an internal AUROC of 0.813 (95% CI: 0.812–0.814) and an external AUROC of 0.868 (95% CI: 0.867–0.868). In the psychiatric group, unspecified dementia (F03) showed an internal AUROC of 0.849 (95% CI: 0.848–0.849) and an external AUROC of 0.862 (95% CI: 0.861–0.863). Discriminative features align with known ECG markers but also provide hints on potentially new markers.
Conclusions: ECG offers significant promise for diagnosing and monitoring neuropsychiatric conditions, with robust predictive performance across internal and external cohorts. Future work should focus on addressing potential confounders, such as therapy-related cardiotoxicity, and expanding the scope of ECG applications, including personalized care and early intervention strategies.
@@
Electrocardiogram abnormalities have been explored as potential biomarkers for neuropsychiatric conditions, but their diagnostic value has not been thoroughly investigated. Recent studies have only begun to examine these abnormalities in exploratory analyses to identify potential correlations.
@@
This study demonstrates that ECG features can effectively predict neuropsychiatric conditions, extending beyond their traditional use as descriptive markers. The findings show high predictive performance across both internal and external datasets. Additionally, the use of an explainable approach allows for a deeper understanding of the correlations between specific ECG markers and neuropsychiatric conditions, confirming existing knowledge while also providing new insights. This research sheds light on the physiological changes that may underlie the interplay between cardiac features and neuropsychiatric disorders.
@@
This work lays the foundation for integrating ECG into the diagnosis and monitoring of neuropsychiatric conditions. It presents a non-invasive, cost-effective tool that could aid in more efficient detection and personalized therapy management in clinical settings, with the potential to transform both research and clinical practices.
@@ 1 Introduction
Neurological and psychiatric conditions are among the most challenging diseases, profoundly affecting both individuals and society. Neurological disorders such as Alzheimer’s disease and Parkinson’s disease lead to progressive cognitive and motor decline, severely limiting patients’ independence and quality of life [1]. Psychiatric conditions like dementia and delirium further complicate mental health, often causing confusion, memory loss, and significant behavioral changes [2]. Notably, psychiatric disorders and neurological conditions are closely interrelated, as they share common neurobiological pathways, genetic factors, and overlapping cognitive, affective, and behavioral symptoms. The distinction between the two is often blurred, complicating diagnosis and treatment, and highlighting the need for integrated, multidisciplinary approaches in neuropsychiatric care [3]. These conditions not only disrupt the lives of those affected but also impose a considerable burden on caregivers and healthcare systems [4]. Early and fast diagnosis is crucial to effectively manage these diseases, but current diagnostic methods can be complex, costly, and sometimes inaccessible [5]. Furthermore, it is important to distinguish Alzheimer’s disease from other major forms of dementia, as dementia describes a group of symptoms, whereas Alzheimer’s disease refers to the underlying pathology [6].
@@
Electrocardiograms (ECG) have long been established as a critical tool in cardiology, used primarily to detect and monitor cardiac conditions by measuring the electrical activity of the heart. ECGs provide vital information about heart rhythms, helping to identify issues such as arrhythmias, ischemia, and other cardiac abnormalities [7]. While ECGs are traditionally associated with cardiac care, recent advancements in medical research have expanded their potential applications. Novel uses of ECG, supported by advancements in data analysis and machine learning, are now being explored in predictive modeling applications across fields beyond cardiology, including the estimation of laboratory values [8] and patient deterioration in emergency departments [9].
The potential link between ECG changes and neurological or psychiatric conditions is an emerging area of study. For instance, conditions like Alzheimer’s and Parkinson’s disease may affect the autonomic nervous system [10], leading to detectable changes in heart rate variability and other ECG patterns. Similarly, psychiatric disorders such as dementia and delirium have been associated with cardiac function [11], potentially reflecting the underlying physiological disruptions. Interestingly, by identifying these subtle ECG changes, one could hope to develop new, non-invasive methods for more efficient detection and monitoring of neurological and psychiatric conditions, offering a promising complement to existing diagnostic tools.
@@ 2.1 Dataset
Figure 1: Diagrammatic illustration of our proposed methodology. The MIMIC-IV-ECG dataset serves as our internal dataset, providing demographic and ECG features used as input for training a tree-based model to diagnose various neuropsychiatric conditions. For external validation, we utilize a second patient cohort from the ECG-View II dataset, extracting the same set of features and neuropsychiatric targets. The conditions are defined based on ICD-10-CM codes.
@@
Female 226,892 (48.50) 375,733 (48.44)
Male 240,837 (51.49) 399,802 (51.55)
@@
Quantile 1 18-53 (23.83) 18-40 (24.03)
Quantile 2 53-66 (25.16) 40-52 (25.75)
Quantile 3 66-78 (25.60) 52-65 (24.94)
Quantile 4 78-101 (25.40) 65-109 (25.28)
@@
Table 1: An overview of the variable characteristics across samples, including demographic information such as gender counts (with ratios) and the median age in years (with interquartile range, IQR), along with the age distribution represented by quantile ratios. It also includes the median (IQR) values for ECG features, such as the RR-interval, PR-interval, QRS-duration, QT-interval, and QTc-interval in milliseconds, as well as the P-wave axis, QRS axis, and T-wave axis in degrees.
Our primary dataset for training and internal evaluation was obtained from the MIMIC-IV-ECG database [12, 13], a subset of a large-scale critical care dataset collected at the Beth Israel Deaconess Medical Center in Boston, Massachusetts. It includes data from patients admitted to the emergency department (ED) and intensive care unit (ICU). Target variables are based on discharge diagnoses encoded using the International Classification of Diseases Clinical Modification (ICD-10-CM). Although a broad range of neoplasm-related codes is explored, this study focuses on those with internal and external validation AUROC scores exceeding 0.7.
To create a comprehensive and harmonized feature set, ECG features from the MIMIC-IV dataset were aligned with those from the ECG-VIEW-II database [14], which serves as our secondary dataset for external validation. ECG-VIEW-II includes data from patients at a South Korean tertiary teaching hospital. The standardized feature set consists of ECG-derived measurements (RR-interval, PR-interval, QRS-duration, QT-interval, QTc-interval in milliseconds; P-wave-axis, QRS-axis, and T-wave-axis in degrees), along with demographic attributes (binary sex and age as a continuous variable).
For the internal dataset, stratified folds were created based on diagnoses, age, and gender distributions, using an 18:1:1 split as described in previous research [15]. A similar stratification procedure was applied to the external dataset to maintain consistency. The training process emphasizes MIMIC-IV-ECG due to its broader ethnic diversity compared to ECG-VIEW-II, which helps improve the model’s generalization across various populations, as shown in prior studies [16], which focused on cardiac conditions, [17] for liver conditions, and [18] for neoplasms. This strategy ensures robust internal training and reliable external validation across ethnically and geographically diverse cohorts.
2.2 Models and evaluation
In this study, we create individual tree-based models using Extreme Gradient Boosting (XGBoost) to tackle binary classification tasks, with a distinct model for each selected ICD-10-CM code. To avoid overfitting, we apply early stopping with a patience of 10 iterations on the validation fold during training. Model performance is assessed using the area under the receiver operating characteristic curve (AUROC) on the validation fold. We report AUROC scores for both the internal test set and an external dataset, along with 95% confidence intervals calculated through empirical bootstrapping with 1000 iterations.
2.3 Explainability
3 Results
3.1 Predictive performance
Code: Description Internal AUROC (95% CI) [Prev.] External AUROC (95% CI) [Prev.]
Neurological conditions
G30: Alzheimer’s 0.8134 (0.8124, 0.8140) [1.04%] 0.8678 (0.8676, 0.8682) [0.13%]
G20: Parkinson’s 0.7796 (0.7790, 0.7799) [0.94%] 0.7897 (0.7900, 0.7905) [0.21%]
G931: Anoxic brain damage 0.7283 (0.7278, 0.7312) [0.5%] 0.7829 (0.7823, 0.7840) [0.05%]
Phychiatric conditions
F01: Vascular dementia 0.7732 (0.7724, 0.7736) [0.57%] 0.8120 (0.8110, 0.8117) [0.11%]
F03: Unspecified dementia 0.8490 (0.8482, 0.8490) [2.6%] 0.8620 (0.8610, 0.8624) [0.12%]
F05: Delirium due to physiological condition 0.7074 (0.7066, 0.7079) [3.41%] 0.7822 (0.7812, 0.7826) [0.12%]
Table 2: Predictive performance results for the neuropsychiatric conditions studied. We present the internal (MIMIC-IV) and external (ECG-View) AUROC performances with 95% confidence intervals, along with the class prevalence of each condition in both datasets.
@@ 3.1 Predictive performance
Table 2 presents the predictive performance of our model for various neuropsychiatric conditions, evaluated using AUROC scores on both internal (MIMIC-IV) and external (ECG-View) test sets. The 95% prediction intervals offer insight into the reliability of these scores. Additionally, the class prevalence for each condition within its respective dataset is provided, offering context on the population distribution. In the MIMIC-IV cohort, prevalences range from 0.5% to 3.41%, while the ECG-View cohort exhibits much lower prevalences, from 0.05% to 0.21%.
Significantly, the conditions with the highest predictive performance in the neurological group include G30: Alzheimer’s, with an internal AUROC of 0.8134 (95% CI: 0.8124–0.8140) and an external AUROC of 0.8678 (95% CI: 0.8676–0.8682). For the psychiatric group, F03: Unspecified dementia achieved an internal AUROC of 0.8490 (95% CI: 0.8482–0.8490) and an external AUROC of 0.8620 (95% CI: 0.8610–0.8624). These findings highlight the reliability of our model in predicting neuropsychiatric conditions, despite differences in prevalence across datasets.
3.2 Explainability
Figure 2: Explainability results for the investigated neurological and phychiatric conditions. The beeswarm plot visualizes through a single dot per feature and sample if the feature contributes positively (right hand side) or negatively (left hand side) to the model prediction. In addition, the color-coding allows to infer if a point is associated with high (red) or low (blue) feature values.
@@ 3.2 Explainability
Figure 2 presents the explainability results based on Shapley values. Across all investigated conditions age is the most important predictive feature, where increased age contributes positively, except for the case of anoxic brain damage. For the investigated neurological conditions, QTc interval is the most important ECG feature for Parkinson’s and anoxic brain damage, where low values and high values of the marker contribute the most positively to each condition respectively. Simiarly, T-wave axis is the second most important ECG feature for the same two conditions, where low values of the marker contributes the most positively to each condition. For the investigated psychiatric conditions, RR-interval is the most important predictive ECG feature across unspecified dementia and delirium due to known physiological conditions, where low values of the marker mainly contribute to the conditions. Similarly, PR-interval is the second most important ECG feature for the same two conditions, where mainly high and low values of the marker contribute positively to the conditions.

### v2

# Explainable and externally validated machine learning for neurocognitive diagnosis via electrocardiograms

@@ Abstract
Background: Electrocardiogram (ECG) analysis has emerged as a promising tool for detecting physiological changes linked to non-cardiac disorders. Given the close connection between cardiovascular and neurocognitive health, ECG abnormalities may be present in individuals with co-occurring neurocognitive conditions. This highlights the potential of ECG as a biomarker to improve detection, therapy monitoring, and risk stratification in patients with neurocognitive disorders, an area that remains underexplored.
Methods: We aim to demonstrate the feasibility to predict neurocognitive disorders from ECG features across diverse patient populations. We utilized ECG features and demographic data to predict neurocognitive disorders defined by ICD-10 codes, focusing on dementia, delirium, and Parkinson’s disease. Internal and external validations were performed using the MIMIC-IV and ECG-View datasets. Predictive performance was assessed using AUROC scores, and Shapley values were used to interpret feature contributions.
Results: Significant predictive performance was observed for disorders within the neurcognitive disorders. Significantly, the disorders with the highest predictive performance is F03: Dementia, with an internal AUROC of 0.848 (95% CI: 0.848-0.848) and an external AUROC of 0.865 (0.864-0.965), followed by G30: Alzheimer’s, with an internal AUROC of 0.809 (95% CI: 0.808-0.810) and an external AUROC of 0.863 (95% CI: 0.863-0.864). Feature importance analysis revealed both known and novel ECG correlates. ECGs hold promise as non-invasive, explainable biomarkers for selected neurocognitive disorders. This study demonstrates robust performance across cohorts and lays the groundwork for future clinical applications, including early detection and personalized monitoring.
1AI4Health Division, Carl von Ossietzky Universität Oldenburg, Oldenburg, Germany.
2Pharmacy Department, Maudsley Hospital, London, UK.
3Department of Psychiatry, University of Oxford, UK.
4Institute of Pharmaceutical Sciences, King’s College London, UK.
5Department of Cardiology, German Heart Center of the Charité-University Medicine, Charité Campus Mitte, Berlin, Germany.
*First author
[…]
@@
Electrocardiogram abnormalities have been explored as potential biomarkers for neurocognitive disorders, but their diagnostic value has not been thoroughly investigated. Recent studies have only begun to examine these abnormalities in exploratory analyses to identify potential associations.
@@
This study demonstrates that ECG features can effectively predict neurocognitive disorders, extending beyond their traditional use as biomarkers for cardiac conditions. The findings show high predictive performance across both internal and external datasets. Additionally, the use of an explainable approach allows for a deeper understanding of the associations between specific ECG features and neurocognitive disorders, confirming existing knowledge while also providing new insights. This research sheds light on the physiological changes that may underlie the interplay between cardiac features and neurocognitive disorders.
@@
This work lays the foundation for integrating ECG into the diagnosis and monitoring of neurocognitive disorders. It presents a non-invasive, cost-effective tool that could aid as a companion tool in more efficient detection and personalized therapy management in clinical settings, with the potential to transform both research and clinical practices.
@@ 1.1 Clinical relevance
Neurocognitive disorders are among the most challenging diseases, profoundly affecting both individuals and society. Disorders such as the underlaying pathology of dementia Alzheimer’s disease [1] and Parkinson’s diseaselead to progressive cognitive and motor decline, severely limiting patients’ independence and quality of life [2]. Other disorders like symphotmatic dementia and delirium further complicate mental health, often causing confusion, memory loss, and significant behavioral changes [3]. Notably, these disorders are closely interrelated, as they share common neurobiological pathways, genetic factors, and overlapping cognitive, affective, and behavioral symptoms. The distinction between them is often difficult, complicating diagnosis and treatment, and highlighting the need for integrated, multidisciplinary approaches in neurocognitive care [4]. These disorders not only disrupt the lives of those affected but also impose a considerable burden on caregivers and healthcare systems [5]. Early and fast diagnosis is crucial to effectively manage these diseases, but current diagnostic methods can be complex, costly, and sometimes inaccessible [6].
1.2 Emerging role of ECG beyond cardiology
@@
Electrocardiograms (ECG) have long been established as a critical tool in cardiology, used primarily to detect and monitor cardiac disorders by measuring the electrical activity of the heart. ECGs provide vital information about heart rhythms, helping to identify issues such as arrhythmias, ischemia, and other cardiac abnormalities [7]. While ECGs are traditionally associated with cardiac care, recent advancements in medical research have expanded their potential applications. Novel uses of ECG, supported by advancements in data analysis and machine learning, are now being explored in predictive modeling applications across fields beyond cardiology, including the estimation of laboratory values [8] and patient deterioration in emergency departments [9].
1.3 Emerging role of neurocardiology
The potential link between ECG abnormalities and neurocognitive disorders is an emerging area of study. For instance, recent studies continue to highlight the complex interplay between mental and cardiovascular health, underscoring the clinical value of early detection and integrated care strategies. For instance, [10] discusses how depression can manifest with cardiovascular symptoms and proposes traditional interventions tailored to the underlying physiological patterns. [11] emphasizes the persistent gap between clinical guidelines and actual practice in monitoring cardiometabolic risk among antipsychotic users, where cardiovascular disease remains a leading cause of death. Additionally, [12] demonstrates that virtual reality-based psychological interventions can effectively reduce anxiety symptoms in acute myocardial infarction patients, offering an example of how digital solutions can be integrated into cardiovascular care. Conditions like Alzheimer’s and Parkinson’s disease may affect the autonomic nervous system [13], leading to detectable changes in heart rate variability and other ECG patterns. Similarly, psychiatric disorders such as dementia and delirium have been associated with cardiac function [14], potentially reflecting the underlying physiological disruptions.
1.4 Current diagnostic practices
To date, few studies have explored the diagnostic utility of ECG for neurocognitive disorders. A retrospective study linked specific ECG features to incident dementia and modestly improved risk prediction, but lacked machine learning to model complex patterns and did not validate findings across independent cohorts, limiting generalizability [15]. A study showed that ECG-derived heart rate variability could detect autonomic dysfunction in patients with REM sleep behavior disorder, an early sign of Parkinson’s disease but was limited by a small sample, lacked broader validation, and did not confirm progression to Parkinson’s [16]. Similarly, a recent study used machine learning to predict neurological outcomes in post-cardiac arrest patients; however, it relied on a single dataset, required EEG for strong performance since ECG-only models underperformed, and relied on synthetic data that may not reflect real-world populations [17].
1.5 Contributions
Based on these observations, this study investigates whether machine learning models trained on ECG data and basic demographic information can accurately identify specific neurocognitive disorders, namely dementia (including Alzheimer’s disease and unspecified dementia), delirium, and Parkinson’s disease. We developed and validated models capable of distinguishing these conditions, demonstrating strong performance across both internal and external datasets. In addition, we provided transparent and explainable insights into the ECG features contributing to each diagnostic category. Our findings highlight the potential of ECG as a clinically useful, low-cost, and non-invasive diagnostic support tool in neurocognitive care, either as a standalone method or, more conservatively, as a complement to existing diagnostic practices. To our knowledge, this is the first study to apply explainable machine learning to ECG data for diagnosing these disorders, with validation across diverse clinical populations confirming its applicability.
@@ 2.1 Dataset
Figure 1: Diagrammatic illustration of our proposed methodology. The MIMIC-IV-ECG dataset serves as our internal dataset, providing demographic and ECG features used as input for training a tree-based model to diagnose various neurocognitive disorders. For external validation, we utilize a second patient cohort from the ECG-View II dataset, extracting the same set of features and neurocognitive targets. The disorders are defined based on ICD-10-CM codes.
Figure 1 contains a schematic representation of our proposed methodology in which the primary dataset for training and internal evaluation was obtained from the Medical Information Mart for Intensive Care (MIMIC)-IV-ECG database [18, 19], a subset of a large-scale critical care dataset collected at the Beth Israel Deaconess Medical Center in Boston, Massachusetts. It includes data from patients admitted to the emergency department (ED) and intensive care unit (ICU). We included patients aged 18 years and older with at least one ECG record and corresponding discharge diagnoses recorded using ICD-10-CM codes between 2008 and 2019. MIMIC-IV-ECG is a publicly available, de-identified dataset, and thus individual patient consent was not required. The use of this dataset was approved by the institutional review boards of the Massachusetts Institute of Technology (MIT) and Beth Israel Deaconess Medical Center through Physionet.
We externally validate our findings using the ECG-VIEW-II database [20]. ECG-VIEW-II includes data from patients at a South Korean tertiary teaching hospital. To create a comprehensive and harmonized feature set, ECG features from the MIMIC-IV dataset were aligned with those from the ECG-VIEW-II database where the standardized feature set consists of ECG-derived measurements (RR-interval, PR-interval, QRS-duration, QT-interval, QTc-interval in milliseconds; P-wave-axis, QRS-axis, and T-wave-axis in degrees), along with demographic attributes (binary sex and age as a continuous variable) which serves as our secondary dataset for external validation. Inclusion criteria for ECG-VIEW-II followed the same logic: adult patients (18+) with at least one ECG record and corresponding ICD-10-CM diagnosis between 1994 and 2015. ECG-VIEW-II is also a publicly available, de-identified dataset released under institutional review board approval, and thus patient consent was not applicable. Target variables are based on discharge diagnoses encoded using the International Classification of Diseases Clinical Modification (ICD-10-CM) [21], we primarly investigate the following disorders; G30: Alzheimer’s disease, G20: Parkinson’s disease, F01: Vascular dementia, F03: Dementia, and F05: Delirium (physiological).
For the internal dataset, stratified folds were created based on diagnoses, age, and gender distributions, using an 18:1:1 split as described in previous research [22]. The external evaluation is performed on the complete cohort. The training process emphasizes MIMIC-IV-ECG due to its broader ethnic diversity compared to ECG-VIEW-II, which helps improve the model’s generalization across various populations, as shown in prior studies [23], which focused on cardiac disorders, [24] for liver disorders, and [25] for neoplasms. This strategy ensures robust internal training and reliable external validation across ethnically and geographically diverse cohorts.
2.2 Models and evaluation
In this study, we create individual tree-based models using Extreme Gradient Boosting (XGBoost) to tackle binary classification tasks, with a distinct model for each selected ICD-10-CM code. To avoid overfitting, we apply early stopping with a patience of 10 iterations on the validation fold during training. We did not apply any imputation into the dataset due to the inherently nature of the utilized model to handle missing data. Model performance is assessed using the area under the AUROC on the validation fold. We report AUROC scores for both the internal test set and an external dataset, along with 95% confidence intervals calculated through empirical bootstrapping with 1000 iterations. To assess model calibration, we present calibration curves based on the internal test set for which we apply model-agnostic calibration and fit isotonic regression models on the validation set and report calibrated test set results. Additionally, we evaluate clinical utility using decision curve analysis, comparing our model’s net benefit to standard strategies such as ”refer all” and ”refer none” [26].
2.3 Explainability
2.4 Study protocol and reporting standards
This study follows a comprehensive protocol aligned with established guidelines, including TRIPOD (Transparent Reporting of a Multivariable Prediction Model for Individual Prognosis or Diagnosis), STARD (Standards for Reporting Diagnostic Accuracy Studies), and MI-CLAIM (Minimum information about clinical artificial intelligence modeling). Detailed documentation is provided in the supplementary material.
3 Results
3.1 Descriptive statistics
@@
Female samples 226,892 (48.50) 375,733 (48.44)
Male samples 240,837 (51.49) 399,802 (51.55)
@@
Quantile 1 age range (%) 18-53 (23.83) 18-40 (24.03)
Quantile 2 age range (%) 53-66 (25.16) 40-52 (25.75)
Quantile 3 age range (%) 66-78 (25.60) 52-65 (24.94)
Quantile 4 age range (%) 78-101 (25.40) 65-109 (25.28)
@@
Table 1: Descriptive statistics of the two ECG datasets, MIMIC-IV-ECG and ECG-View II. The table summarizes demographic variables gender (absolute sample count and percentage) and age (median and interquartile range) along with age group distributions by quantiles. ECG features are reported as median (IQR), with temporal intervals ($R$, PR, QRS, QT, QTc) given in milliseconds (ms), and electrical axes (P-wave, QRS, T-wave) in degrees ().
Table 1 highlights key demographic and ECG feature differences between MIMIC-IV-ECG and ECG-View II. Both datasets have a nearly identical gender distribution with slightly more male than female samples. However, MIMIC-IV-ECG includes an older population (median age: 66 vs. 52 years), which may influence ECG characteristics. Notable differences include a shorter RR-interval in MIMIC-IV-ECG (769 ms vs. 857 ms) and a higher QTc-interval (447 ms vs. 421 ms), suggesting potential variations in heart rate and repolarization patterns. Additionally, the QRS axis shows a substantial shift (13° vs. 48°), which may reflect the differences in patient populations. A detailed comparison of ECG features between diagnostic-positive and -negative groups across the MIMIC-IV and ECG-VIEW II datasets is presented in the supplementary Table 3. The analysis reveals modest yet consistent feature differences, suggesting potential disease-related alterations in ECG signals in an exploratory manner.
3.2 Discriminative, calibration, decision, and interpretability analysis
Figure 2: Model evaluation results across five neurocognitive disorders. Columns represent: (1) Discriminative performance (AUROC with 95% CI), (2) Calibration plots, (3) Net benefit decision curves, and (4) SHAP-based feature importance. Rows correspond to the following ICD-10 codes; G30: Alzheimer’s disease, G20: Parkinson’s disease, F01: Vascular dementia, F03: Unspecified dementia, and F05: Delirium due to physiological condition.
Comprehensively outlined by [28], assessing the predictive performance of AI algorithms in medicine has to cover several different categories, most notably, discrimination, clinical utility. We assess discimination in terms of receiver operating curves/AUROC scores, calibration through calibration plots and clinical utility through decision curve analysis/net benefit [26]. This is complemented through an explainability analysis by means of SHAP values. We showcase each of the four categories as columns in Figure 2.
@@ 3.2 Discriminative, calibration, decision, and interpretability analysis
The first column in Figure 2 presents the predictive performance of our model for various neurocognitive disorders, evaluated using AUROC scores on both internal (MIMIC-IV) and external (ECG-View) test sets. The 95% prediction intervals offer insight into the reliability of these scores. Additionally, the class prevalence for each condition within its respective dataset is provided, offering context on the population distribution and demostrating predictive performance robustness due to their prevalence missmatch. In the MIMIC-IV cohort, prevalences range from 0.5% to 3.41%, while the ECG-View cohort exhibits much lower prevalences, from 0.05% to 0.21%. Significantly, the disorders with the highest predictive performance in the is F03: Dementia, with an internal AUROC of 0.848 (95% CI: 0.848–0.848) and an external AUROC of 0.865 (0.864-0.965), followed by G30: Alzheimer’s, with an internal AUROC of 0.809 (95% CI: 0.808–0.810) and an external AUROC of 0.863 (95% CI: 0.863–0.864). These findings highlight the reliability of our model in predicting neurocognitive disorders, despite differences in prevalence across datasets.
The second column in Figure 2 illustrates the calibration of the model predictions for each condition, comparing predicted probabilities with observed outcomes. Across all conditions, our models demonstrate good calibration, indicating that the predicted probabilities are consistent with actual event rates. This is particularly relevant in clinical applications, where over- or underestimation of risk could have critical consequences. The reliability of these probability estimates further supports the model’s utility in risk stratification and decision support.
The third column in Figure 2 presents decision curve analyses, assessing the clinical utility of our models. For all examined disorders, the models provide a higher net benefit across a range of threshold probabilities compared to the ’refer-all’ and ’refer-none’ strategies. This suggests that using our models to guide referral decisions could lead to better patient outcomes while minimizing unnecessary evaluations. The consistent superiority of our models in net benefit reinforces their practical value in aiding early detection and triage for neurocognitive disorders.
@@
The fourth column in Figure 2 represents the result of the interpretability analysis. Across all disorders investigated, age is the most important predictive feature, where increased age contributes positively. For the investigated neurocognitive disorders, QTc interval is the most important ECG feature for Parkinson’s, where low values and high values of the marker contribute the most positively to each condition respectively. Similarly, T-wave axis is the second most important ECG feature for the same two disorders, where low values of the marker contributes the most positively to each condition. For the investigated psychiatric disorders, RR-interval is the most important predictive ECG feature across unspecified dementia and delirium due to known physiological disorders, where low values of the marker mainly contribute to the disorders. Similarly, PR-interval is the second most important ECG feature for the same two disorders, where mainly high and low values of the marker contribute positively to the disorders.
Finally, we investigate ICD-10 codes with strong internal discriminative performance on the MIMIC-IV-ECG dataset that could not be extermally evaluated on the ECG-View II. Some of the high-performing categories include eating disorders such as F502: Bulimia nervosa (AUROC = 0.953) and F509: Eating disorder, unspecified (0.916), personality disorders such as F602: Antisocial personality disorder (0.952), and conduct disorders such as F91: Conduct disorders (0.878). Additionally, several substance-related disorders showed strong internal performance, including F1510: Other stimulant abuse, uncomplicated (0.903), F1220: Cannabis dependence, uncomplicated (0.891), F1310: Sedative, hypnotic or anxiolytic abuse, uncomplicated (0.884), F1910: Other psychoactive substance abuse, uncomplicated (0.895), and F1012: Alcohol abuse with intoxication (0.873). The full list of these codes is provided in the supplementary Table LABEL:tab:noext, including 8 codes with AUROCs above 0.9 and 40 codes between 0.8 and 0.9, representing promising directions for future investigation. Additionally, the supplementary Table 2 lists conditions that showed good performance on the internal test set but failed to generalize externally. For instance, G931 (Anoxic brain damage) was excluded from the final analysis despite strong internal and external AUROC values, due to poor calibration and limited clinical utility (see the supplementary Figure 1).